What are the roots of (4x-3)^2=36?
Answer by MaxLiebling(13)

(4x-3)^2=36
4x-3 = +/-6
4x = +/-6 + 3
4x = -3 or +9
x = -3/4 or +9/4
proof:
x1 = -3/4
(4x-3)^2 = (4*(-3/4)-3)^2 = (-3-3)^2 = -6^2 = 36 ok.
x2 = +9/4
(4x-3)^2 = (4*(+9/4)-3)^2 = (+9-3)^2 = 6^2 =36 ok.